Given the function g described by g(x)=6x find each of the folowing.
a.g(-10)
b.g(13)
c.g(28)

Found 2 solutions by user_dude2008, neatmath:
Answer by user_dude2008(1862)

a.g(-10) = 6(-10) = -60
b.g(13) = 6(13) = 78
c.g(28) = 6(28) = 168

Answer by neatmath(302)

To solve these kinds of problems, every time you see an x, replace it with the number in the parentheses!

We are just evaluating functions at given points.

In other words, we are given the x value, and we want to find the g(x) value, or y value!

For example, we have g(x)=6x

If we want g(1) we would just say that g(1)=6*1

Therefore g(1)=6

So, for your problems,

a) g(-10)=6*(-10)=-60

b) g(13)=6*13=78

c) g(28)=6*28=168

I hope this helps!